OpenStudy (anonymous):

) Solve for x: 2x - 4 = -3

OpenStudy (idealist10):

2x-4=-3 2x=-3+4 2x=1 x=1/2 Answer: 1/2

Nnesha (nnesha):

4 is subtracting by 2x so what is opposite of subtract ???

Nnesha (nnesha):

opposite of subtract is add so first add 4 both side

Nnesha (nnesha):

and then on left side you have left 2x so 2 is multiplying by x what is opposite of multiplying ???

Nnesha (nnesha):

isn't that dividing ?? so divide both side by 2 then you get your answer

OpenStudy (idealist10):

Do you get it, @kamontel1?